Use differentials to approximate the volume of material needed to make a rubber ball if the



radius of the hollow inner core is 2 đť‘–đť‘›., and the thickness of the rubber is 1



8



đť‘–đť‘›.

1
Expert's answer
2021-11-29T19:10:57-0500

Volume of the sphere is "V=\frac{4}{3}\pi r^3"

"dV=\frac{4}{3}\pi (3r^2)dr"

"=4 \pi r^2dr\\=4\pi (4)^2(0.2)=12.8 \pi"